OpenStudy (anonymous):

How do i do this? Which radical expression is equivalent to (3^1/4) ^6

OpenStudy (lucaz):

radical expression is when you have a radical, or root, on it

OpenStudy (ranga):

(a^m)^n = a^mn Then write the fractional exponent as a radical.

OpenStudy (anonymous):

3 ^ 1/4/ 6?

OpenStudy (lucaz):

it's 3^(1/4) ?

OpenStudy (anonymous):

or i could multiply them

OpenStudy (lucaz):

as ranga said, you multiply the exponents (1/4)*6

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

(a^m)^n = a^(mn) or Multiply exponents to get 3^6/4 or 3^3/2 This equals the square root of 3^3 or the sq rt of 27. This simplified is 3 times the sq rt of 3

OpenStudy (anonymous):

3 ^ 3/2 = 2 sqrt (3^3) ?

OpenStudy (ranga):

You multiply the exponents. (3^1/4) ^6 = 3^(1/4 * 6) = 3^(6/4) = 3^(3/2) = sqrt(3^3) = sqrt(27)

OpenStudy (anonymous):

wait in the lesson they always put down 2 under "m" when there was no "n" why?

OpenStudy (ranga):

not sure I understand your last question.

OpenStudy (anonymous):

if i have sqrt5 ^ 4 when i rewrite it tit would be 5^ 4/2

OpenStudy (anonymous):

@ranga is smart take his advice he is right!

OpenStudy (anonymous):

ok :)

OpenStudy (ranga):

Yes. Taking a square root is SAME AS having an exponent of 1/2. sqrt(A) = A^(1/2)

OpenStudy (anonymous):

oh ok, on this one tell me if im right i would have: 4/3 x 1/2 = 2/3 2sqrt.5^3 or 5^2/3?

OpenStudy (ranga):

Not 4/3. It should be 3/4. Taking the 4th root is same as exponent 1/4 Taking the nth root is same as raising it to ^1/n

OpenStudy (anonymous):

m/n = 3/4 right so 3/4 x 1/2 = 3/8 so 5^3/8

OpenStudy (anonymous):

3/4+1/2=1 and 1/4

OpenStudy (anonymous):

does that say 5 times 1 and 1/4

OpenStudy (anonymous):

if so then= 5 and 1/4

OpenStudy (ranga):

Inside the radical is 5^3. Outside the radical is that little 4 which means the 4th root. 4th root of (5^3) = ( 5^3 )^1/4 = 5^(3/4) which is raised to 1/2 ( 5^(3/4) )^1/2 = 5^(3/4 * 1/2) = 5^(3/8)

OpenStudy (anonymous):

because (4sqrt.5^3) is in the ( ) that means that (5^3/4) would be to so like the last if im correct i would multiply the exponents then simplify if possible for my answer right?

OpenStudy (anonymous):

if you need help with fractions always message me i can help any kind of fraction

OpenStudy (ranga):

yes. But be careful when writing math expression without the math editor. 4sqrt 5^3 will be confused with 4 * sqrt(5^3) and that is not correct. This is the 4th root of 5^3.

OpenStudy (anonymous):

ok :) i have two more that i need to be check i think i am starting to get this on this i would end up with 16^3/9 3/9 = 1/3 so i would have 3rd root of 16 ^1 thus the answer would be 3rd root of 16 right?